If a specific amount water A at 50 degree Celsius is mixed with water B at 90 degree Celsius that is 40 grams less than
water A the resulting temperature is 65.33 degree Celsius. What is the mass of water A and water B? Assume
no heat is lost to the surrounding.
1
Expert's answer
2020-11-09T14:08:26-0500

ma cΔT= mb cΔT

Let mass of water A and B be a and b respectively

a=b+40 (given in the question)

(b+40)×c×(65.33-50)= b ×c× (90-65.33)

(b+40)×15.33= b×24.67

15.33b+ 613.2= 24.67b

613.2=24.67b-15.33b

613.2 = 9.34b

b=65.65g


a=b+40

a= 65.65+40

a= 105.65g

Mass of water A and water B is 105.65g and 65.65g respectively